2. Interest Rate Parity: Suppose the US dollar interest rate and the Canadian dollà (a) Suppose the expected future exchange rate (BeADIUSp) is fixed at USD 1.05 for 1 CAD, (b) Suppose ECAD/USD remains the same while the US interest rate decreases to 8%. If (RusD) interest rate (RCA) are the same, 10% per year what is the equilibrium current exchange rate, ECADIUSD the Canadian interest rate remains the same as before, what is the new equilibrium ECAD/USD?
Answer #1

Formally interest rate parity can be written as

\large \frac{1+i_\$}{1+i_{CAD}} = \frac{F_{\$/CAD}}{S_{\$/CAD}}

US interest rate = 10%

Canadian interest rate = 10%

Future exchange rate = $ 1.05 per CAD

\large \frac{1+0.1}{1+0.1} = \frac{\$ 1.05 /CAD}{S_{\$/CAD}}

\large 1 = \frac{\$ 1.05 /CAD}{S_{\$/CAD}}

\large S_{\$/CAD} = \$ 1.05 /CAD

Equilibrium exchange rate = $ 1.05 / CAD

B. US interest rate = 8%

\large \frac{1+0.08}{1+0.1} = \frac{\$ 1.05 /CAD}{S_{\$/CAD}}

\large \frac{1.08}{1.1} = \frac{\$ 1.05 /CAD}{S_{\$/CAD}}

\large S = \$ 1.05 /CAD *\frac{1.1}{1.08}

S = $ 1.0694 / CAD

Please contact if having any query thank you. Will be obliged if you contact before rating up. Thank you.